Circular No. 6809

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1997ef, 1997ei, 1997fb, 1997fc, 1998A A. V. Filippenko and E. C. Moran, University of California at Berkeley,

report the results of spectroscopy done on Jan. 17 with the Keck-2 telescope: “SN 1997fb and SN 1997fc (cf. IAUC 6807) are type-Ia supernovae within ∼ 2 weeks past maximum brightness, although the spectrum of the latter appears somewhat peculiar (perhaps due to starlight contamination). Their redshifts are consistent with that of Abell 3301 (0.054). SN 1998A (cf. IAUC 6805, but 40′′ west, not east, of the galaxy nucleus) is a type-II supernova with well-developed P-Cyg Balmer lines. SN 1997ei is of type Ic (cf. IAUC 6802), not type Ia (cf. IAUC 6796, 6800). A near-infrared spectrum of SN 1997ef (cf. IAUC 6786, 6798) reveals strong, broad Ca II 860-nm emission, but this a characteristic of many supernova subtypes; the exact classification of SN 1997ef is still uncertain.”

SUPERNOVA 1998B IN ANONYMOUS GALAXY J. Mueller reports her discovery of a supernova (mag ∼ 18.5) on plates

taken on Jan. 1 and 2 with the 1.2-m Oschin Schmidt Telescope in the course of the second Palomar Sky Survey. SN 1998B is located 7′′ east and 1′′.5 south of the center of a galaxy at α = 7^{h}46^{m}02^{s}.1, δ = +18^{o}43′02′′ (equinox 2000.0). There is no object at this position on the Digital Sky Survey. Filippenko and Moran report that a Keck-telescope spectrum of the object, obtained as above, shows that it is a type-Ia supernova, 3–4 weeks past maximum brightness. The redshift of the host galaxy is 0.045.

COMET C/1997 T1 (UTSUNOMIYA)

Ephemeris from orbital elements on MPC 31069:

1998 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Jan. 17 18^{h}49.^{m}71 + 0^{o}54.^{′}8 2.275 1.468 27.^{o}0 17.^{o}7 10.5
27 18 49.88 − 0 48.1 2.275 1.528  31.7 19.8 10.7
Feb. 6 18 49.07 − 2 22.3 2.239 1.598  39.0 22.8 10.8
16 18 46.79 − 3 52.2 2.170 1.676  47.9 25.9 10.9
26 18 42.39 − 5 22.5 2.074 1.761  57.9 28.5 10.9
Mar. 8 18 35.02 − 6 57.9 1.956 1.851  69.1 30.1 11.0
18 18 23.58 − 8 43.5 1.824 1.945  81.5 30.4 11.0
28 18 06.69 −10 43.7 1.691 2.042  95.3 29.1 11.0
Apr. 7 17 42.82 −12 59.9 1.569 2.141 110.9 25.9 11.0
17 17 10.95 −15 26.3 1.477 2.242 128.3 20.6 11.0
1998 January 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green